# Finite Automata Extraction: Low-data World Model Learning as Programs from Gameplay Video

### v1

@@ Abstract
World models are defined as a compressed spatial and temporal learned representation of an environment. The learned representation is typically a neural network, making transfer of the learned environment dynamics and explainability a challenge. In this paper, we propose an approach, Finite Automata Extraction (FAE), that learns a neuro-symbolic world model from gameplay video represented as programs in a novel domain-specific language (DSL): Retro Coder. Compared to prior world model approaches, FAE learns a more precise model of the environment and more general code than prior DSL-based approaches.
Introduction
World models are defined as a compressed spatial and temporal learned representation of an environment (Ha and Schmidhuber 2018a). They are predictive models that mimic the environment and approximate future states based on the actions of an agent. Researchers typically apply World model learning to reinforcement learning or other automated game playing approaches.
Recently, some researchers have moved towards motivating world model learning for game creation (Parker-Holder et al. 2024; Valevski et al. 2025). However, there are some issues with this approach. Since prior instances of these playable world models are all neural networks, they are a poor fit for game developers. They cannot make changes to the behavior of the model as they could with traditional game code. Training these models also requires a large amount of data and computation, resources which are not available to the majority of developers. If we had a world model learning approach that better suited the needs of game developers in terms of an editable code representation and only requiring low data, it could be a powerful tool for development and analysis.
@@ Introduction
This paper presents a novel neuro-symbolic approach to world mode learning from gameplay videos. Our pipeline learns a dictionary of sprites in the video and then searches over Retro Coder, a novel Domain Specific Language to extract programs for each of the learned sprites. This approach would give developers access to the underlying code behind the world model. Our approach also uses substantially less data and is less computationally intensive than previous world model approaches.
@@
We apply our approach on two domains, Pac-Man and River Raid. We then compare our results to GameGan, a neural world model (Kim et al. 2020), and Game Engine Learning, a DSL-based world model (Guzdial, Li, and Riedl 2017a). In this paper, we present the following contributions:
@@
  • A novel Domain Specific Language, Retro Coder.
@@
Related Work
In this section we first cover prior work in program synthesis, given that Finite Automata Extraction (FAE) relies on it to learn appropriate programs in our Domain Specific Language, Retro Coder. We then cover prior work in world models.
Program Synthesis
@@
Program synthesis refers to the task of generating programs that satisfy user needs, commonly expressed via a set of constraints or heuristic (Gulwani et al. 2017). Commonly, developers define a search space of programs from which a desired program can be selected. Program synthesis has long been an area of interest for artificial intelligence research (Manna and Waldinger 1975; Waldinger and Lee 1969; Pnueli and Rosner 1989) with more recent efforts leveraging neural network-based approaches (Balog et al. 2017; Devlin et al. 2017; Ellis et al. 2018, 2023).
@@
Within the specific context of games, program synthesis has been used for generating boss fights (Butler, Siu, and Zook 2017), solving grid-based games (Silver et al. 2020), generating strategies for RTS games (Marino et al. 2021) and generating goal-based minigames (Davidson et al. 2025). More commonly program synthesis is used for automated control. (Gu et al. 2024) use program synthesis for representing policies for agents in the Overcooked environment. Similarly (Eberhardinger, Maucher, and Maghsudi 2023) synthesize programs to explain an agent’s decision-making process in the Atari environment. More recent work has also looked at defining a framework for leveraging large language models to assist with synthesizing programs for accomplishing a variety of in-game tasks (Eberhardinger et al. 2024). The majority of this prior work does not focus on program synthesis for world model learning, we address exceptions in the next subsection.
World Models
@@
Figure 1: System Overview of Finite Automate Extraction (FAE). We first extract frames from gameplay video, then learn a neural sprite dictionary and a symbolic representation of those frames, and then learn programs for each of the sprite in the sprite dictionary
@@
World models are most commonly used for training reinforcement learning agents (Ha and Schmidhuber 2018b; Bamford and Lucas 2020; Hafner et al. 2020, 2023). A recent focus has been in training world models that enable real-time interaction via user inputs (Menapace et al. 2021; Wu et al. 2024). While some of these works (Bruce et al. 2024; Valevski et al. 2025; Parker-Holder et al. 2024; Yu et al. 2025) are often presented as models capable of mimicking or generating games, they typically operate by generating video frames conditioned on user input in real-time and lack explicit knowledge of underlying game rules and mechanics. Due to their size, the majority of the recent world models cannot run in real-time on consumer hardware and are not open source. For these reasons we compare to only GameGan in our experiments (Kim et al. 2020).
@@
Some prior work has looked at more explicitly modeling a game mechanics such as Game Engine Learning, which employed a Domain Specific Language of if-then rules in a program synthesis-based symbolic learning approach (Guzdial, Li, and Riedl 2017a). We compare against this approach as well in our experiments as the most similar example of purely symbolic prior work. There are other symbolic approaches, such as learning hybrid automata (Summerville et al. 2017) but they depend on access to an emulator instead of purely gameplay video.
System Overview
@@
In this section, we cover our proposed neurosymbolic approach to learn programs from gameplay video: Finite Automata Extraction (FAE). We give an overview of our pipeline in Figure 2. We begin by splitting the gameplay video data into RGB frames. We then train a self supervised model based on the Marionette architecture (Smirnov et al. 2021) on these frames to extract a neural sprite dictionary, a bounded set of sprites. After training the model, we extract a learned symbolic representation of the frames. We then learn a program for each learned sprite to replicate their behavior in the gameplay video. We learn each program by searching over Retro Coder, a novel domain-specific language.
Representation Extraction
@@
We require a representation sprites and the positional information of each sprite in each frame for the later program learning step. We decided to proceed with a grid representation, where each grid is a matrix of symbolic values, where each value corresponds to one learned sprite. This enables us to handle collisions discretely. In order to obtain the grid representation, we chose to learn the sprite dictionary consisting of $l$ number of sprites using a variant of the Marionette architecture (Smirnov et al. 2021). The original Marionette is a regularized Variational Autoencoder (VAE) that learns to extract a discrete set of game sprites from video for the purposes of video reconstruction.
@@
We modified the original Marionette network to add an option to specify the background image, which is helpful for domains with a constant background. We changed the model to subtract the background image from both the input and output to ensure the model optimized only the sprites and reconstruction. The original Marionette has multiple different loss components, including one for frame reconstruction. We added an additional multiplicative weight ($\gamma$) to this loss to further bias the model towards better reconstructions, which indirectly incentivizes learned sprite quality. We set $\gamma=10$ based on early experimentation.
@@
After training, for all the RGB frames $F_{1},...F_{n}$}, we extract their grid representations $G_{1},...G_{n}$}. Marionette employs a latent representation of frames in terms of grids of sprite embeddings. Thus we could extract these grids via passing the frames through the model and using the softmax weights from the model to determine the sprite present in each grid cell. This resulted in the $w\times h$ grid where each cell referenced a sprite in the sprite dictionary.
Retro Coder
@@
We wanted to define a DSL that would be able to make the behavior of the sprites learnable. We thus defined our custom Domain Specific Language, Retro Coder. We learn a program for each sprite individually. This is done for the ease of potential game developers who might use this pipeline, since they be able to make changes to the behavior of a particular sprite easily this way. Each program for each sprite consists of if-then statements as this modular structure allows us to easily extract information about one entity. Thus Retro Coder consists of condition functions (if) and action functions (then). Need to say here how conditions are structured, is it one condition per if? etc. Since this pipeline is made for 2D games, we focused on common sprite behaviors and conditions for these functions. These consist of movement logic, entity transformation logic, and entity collision logic. The condition functions are:
  • exists_in_map(entity)exists\_in\_map(entity) checks if an object of type entityentity exists in the frame
@@
  • $neighboring(entity)$ checks if the object of the program itself is next to an object of type entityentity,
@@
  • $neighbours(entity1,entity2)$ checks if an object of type $entity1$ is next to an object of type $entity2$. Since we are using a grid representation where one sprite can exist at one location in a moment of time, we treat neighboring sprites as collisions.
@@
  • follow_entity(entity)follow\_entity(entity) moves the object toward the nearest object of type entityentity
  • follow_direction(direction)follow\_direction(direction) moves the object toward a direction (Up, down, left, right).
  • change_to_entity(entity)change\_to\_entity(entity) changes the object’s type to entityentity
  • follow_target_location(x,y)follow\_target\_location(x,y), which moves the object towards to location $x,y$.
Finite Automata Extraction
@@
 $bestdistance\leftarrow getdistancetonextframes$ for $i\leftarrow 1$ to $len(listofIndices)$ do
    $bestdistance\leftarrow 0$;
    for $j\leftarrow 1$ to batchSizebatchSize do
       $frame\leftarrow listofIndices[i+j]$;
       $bestdistance\leftarrow bestdistance+loss(grids[frame+1],grids[frame])$;
      end for
@@
       $bestNeigbour\leftarrow None$;
@@
          $newdistance\leftarrow 0$;
          for $j\leftarrow 1$ to batchSizebatchSize do
             $frame\leftarrow listofIndices[i+j]$;
             $current\leftarrow grids[frame]$;
             $next\leftarrow grids[frame+1]$;
             $nextPrediction\leftarrow applyProgram(current,neighbour)$;
             $newdistance\leftarrow newdistance+distance(nextPrediction,next)$;
            end for
          if $newdistance<bestdistance$ then
             $newdistance\leftarrow bestdistance$;
@@
             $distanceDecreased\leftarrow True$
@@
 $programs.\texttt{append}(program)$
@@
Algorithm 1 Pseudocode for Finite Automata Extraction
@@
This section covers the program synthesis part of the pipeline. From the sprite dictionary, we hand-identify sprites to learn programs. This might be based on domain knowledge but can also be general at times. For example, the sprite that is controlled by the player would typically not be learned. We then learn a program for each of these hand-identified sprites with Finite Automata Extraction, as seen in 1. We learn one program at a time, as this greatly reduces the search space. For a sprite $s$, we first scan $G_{1},...G_{n}$} to get a list of indices $I_{1},...I_{s}$} where $G_{I_{i}}$ contain $s$. We then go through the indices sequentially, for every $i$, compute the distance $D=\sum_{x=i}^{i+b}A.E(G_{I_{x+1}},G_{I_{x}})$ where $A.E$ is the Absolute Error and $b$ is the batch size. We set $b=3$ based on experimentation. We then search for neighbour programs to minimize $D$. The neighbour programs can either have another if-then statement, or remove an if-then statement from the program.
Evaluation
We evaluate our model from two perspectives: the quality of the generated programs, which assesses the backend of the model, and the quality of the world model, which assesses the front end. We test our model on two domains, Pac-Man and River Raid. The two domains vary in mechanics and thus serve as a good representation of retro 2D games.

### v2

# Finite Automata Extraction: Low-data World Models as Programs from Gameplay Video

@@ Abstract.
World models are often neural network-based models that attempt to approximate an entire video game. Existing world models are not practical for game developers due to their large scale, lack of accessibility, their prompt-based interfaces, and their nature as black box models where developers cannot access the code. In this paper, we propose an approach, Finite Automata Extraction (FAE), that learns a neuro-symbolic world model represented as programs in a novel domain-specific language (DSL): Retro Coder. The neuro-symbolic world model structure allows for low data learning, meaning we can learn these models from a single gameplay video. This model’s structure means that developers can directly access code to change the behaviour of the model, and it’s low data cost makes it more accessible. In a comparison to prior neural world model approaches, FAE learns a more precise model of the environment and it learns more general code than prior DSL-based approaches.
World Models, Video Games, Neuro-Symbolic, Machine Learning, Artificial Intelligence
copyright: acmlicensedjournalyear: 2026doi: 10.1145/3815598.3815632conference: Foundations of Digital Games; August 10–13, 2026; Copenhagen, Denmarkbooktitle: Foundations of Digital Games (FDG ’26), August 10–13, 2026, Copenhagen, Denmarkisbn: 979-8-4007-2495-4/2026/08ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Computer visionccs: Computing methodologies Machine learningccs: Computing methodologies Semi-supervised learning settings
1. Introduction
Modern video games are primarily made using game engines(Sharma et al., 2024). Game developers are typically required to have technical depth and breadth in order to use these engines, needing knowledge of aspects such as programming languages, mathematics, physics concepts, AI implementations and so on. As a result, the barrier to entry to becoming a game developer can be quite high. With the arrival of the third AI summer, there has been a trend to develop tools that can lower this barrier (Ternar et al., 2025). Some industry researchers have recently proposed world models as a tool for game development (Valevski et al., 2025).
World models are a compressed spatial and temporal learned representation of an environment (Ha and Schmidhuber, 2018). They are predictive models that mimic an environment and approximate future states based on the actions of an agent. For example, creating a simplified, playable clone of Pac-Man in a neural network as in NVIDIA’s GameGAN (Valevski et al., 2025). Researchers typically apply world model learning to reinforcement learning or other automated game playing approaches (Ha and Schmidhuber, 2018). Recently, some industry researchers have moved towards motivating world model learning for game creation (Parker-Holder et al., 2024; Valevski et al., 2025). They claim that world models can act as game engines, and a particular subset of the latent space/world model can be devoted to a single playable game. Thus game development would become as simple as searching through this latent space of possible games. However, there are some issues with this approach. These models do not have a game state or memory, and thus have a tendency to hallucinate or lose track of offscreen objects (Suresh et al., 2025). This could make it hard for developers to make a cohesive gameplay loop, since the mechanics in the environment would be unpredictable and unreliable. Since these playable world models are all neural networks, they are also a black box to the developer (Buhrmester et al., 2019). They cannot make changes to the behaviour of the model as they could with traditional game code. Training these models also requires a large amount of data and computation, resources which may not be available to the majority of developers. As a result, the only viable option for developers would be to use pretrained models, which further reduces their autonomy. Such models like Genie 3, can be accessed with a paid subscription (Ball et al., 2025). Recently, OpenAI shut down SORA, a video generation model, raising concerns around long-term access.
@@ 1. Introduction
If we had a world model learning approach that better suited the needs of game developers in terms of an editable code representation and only requiring low data, it could be a powerful tool for development. This paper presents Finite Automata Extraction (FAE): a novel neuro-symbolic approach to world model learning from a single gameplay video. A neuro-symbolic pipeline uses a combination of neural network architectures and symbolic approaches. Our pipeline learns a dictionary of sprites in a gameplay video using a neural network and then searches over Retro Coder, a novel Domain Specific Language (DSL) to extract symbolic programs for each of the learned sprites. A DSL is a custom simplified language created for a specific domain. This approach would give developers access to the underlying code behind the world model. This would remove the black box nature of a traditional world model, allowing developers to directly alter the mechanics of the game and remove unintended unpredictability. Our approach also uses substantially less data and is less computationally intensive than previous world model approaches. We argue that this would make it more accessible to more game developers. As an initial implementation, our pipeline targets simple grid based 2D games. In this paper we present the idea and initial evidence to support the feasibility of our pipeline.
@@
We apply our approach on two domains, Pac-Man and River Raid and compare our results to GameGAN, a neural world model (Kim et al., 2020), and Game Engine Learning, a DSL-based world model (Guzdial et al., 2017b). In this paper, we present the following contributions:
@@
  • A novel Domain Specific Language: Retro Coder.
@@
2. Background
In this section we first cover the technical background knowledge needed to understand our work.
2.1. Program Synthesis
@@
Program synthesis refers to the task of generating programs that satisfy user needs (Gulwani et al., 2017b). Commonly, developers define a search space of programs from which a desired program can be selected. Typically, this is done by selecting or writing a Domain Specific Language, as the search space in a custom DSL is usually much smaller than a Turing complete language. The developers then define a set of constraints or heuristic for the desired program (Gulwani et al., 2017a). For instance, they might give a series of input output pairs for the program. The goal of the search algorithm then would be to search for a program that satisfies most, if not all of these examples. Program synthesis has long been an area of interest for artificial intelligence research (Manna and Waldinger, 1975; Waldinger and Lee, 1969; Pnueli and Rosner, 1989) with more recent efforts leveraging neural network-based approaches (Balog et al., 2017; Devlin et al., 2017; Ellis et al., 2018, 2023).
2.2. MarioNette
MarioNette is a self-supervised neural network model that decomposes video frames into sprites, i.e distinct pieces of image data stored in a game(Smirnov et al., 2021). It is a regularized Variational Auto-encoder that takes in video frames as input and reconstructs them. It does this by learning a sprite dictionary, a bounded set of neural sprites, i.e., sprite embedding learned by Marionette approximating true sprites. It uses this dictionary to reconstruct the frame. To do this, it employs a grid representation of the frame, where each grid cell contains a sprite from the sprite dictionary. The model has been found successful at this task for the domain of Super Mario Bros.. We leverage Marionette as the basis for the neural network in our approach. By learning sprites, we can avoid the requirements of prior DSL-based world models in having to provide all possible sprites as the developer (Guzdial et al., 2017b).
3. Related Work
In this section we first cover prior work in program synthesis, given that Finite Automata Extraction (FAE) relies on it to learn appropriate programs in our Domain Specific Language, Retro Coder. We then cover prior work on world models.
3.1. Program Synthesis in Games
@@
Within the specific context of games, program synthesis has been used for generating boss fights (Butler et al., 2017), solving grid-based games (Silver et al., 2020), generating strategies for RTS games (Marino et al., 2021) and generating goal-based minigames (Davidson et al., 2025). More commonly program synthesis is used for automated control. Gu et al. use program synthesis for representing policies for agents in the Overcooked environment (Gu et al., 2024). Similarly Eberhardinger et al. synthesize programs to explain an agent’s decision-making process in the Atari environment (Eberhardinger et al., 2023). More recent work has also looked at defining a framework for leveraging large language models to assist with synthesizing programs for accomplishing a variety of in-game tasks (Eberhardinger et al., 2024). The majority of this prior work does not focus on program synthesis for world model learning, we address exceptions in the next subsection.
3.2. World Models
@@
Figure 1. System Overview of Finite Automate Extraction (FAE). We first extract frames from gameplay video, then learn a neural sprite dictionary and a symbolic representation of those frames, and then learn programs for each of the sprites in the sprite dictionary
@@
A recent focus has been in training world models that enable real-time interaction via user inputs (Menapace et al., 2021; Wu et al., 2024). While some of these works (Bruce et al., 2024; Valevski et al., 2025; Parker-Holder et al., 2024; Yu et al., 2025; Ball et al., 2025) are often presented as models capable of mimicking or generating games, they operate by generating video frames conditioned on user input in real-time and lack explicit knowledge of underlying game rules and mechanics. Due to their size, the majority of the recent world models cannot run in real-time on consumer hardware and are not open source. For these reasons we compare to only GameGAN in our experiments (Kim et al., 2020). GameGAN uses a Generative Adveserial Network to predict the next frame, It takes in the previous frame and keyboard actions as its input.
@@
Some prior work has looked at more explicitly modelling of game mechanics such as Game Engine Learning, which employed a Domain Specific Language of if-then rules in a program synthesis-based symbolic learning approach (Guzdial et al., 2017b). Our approach differs from this in two ways. Firstly, our pipeline includes learning the sprites, which is not the case with Game Engine learning. Secondly, we use a bottom up approach to learn the programs, as opposed to Game Engine Learning, which uses a top down approach. As a result, our approach learns rules that are more general. We compare against this approach as well in our experiments as the most similar example of purely symbolic prior work. There are other symbolic approaches, such as learning hybrid automata (Summerville et al., 2017) but they depend on access to an emulator instead of purely gameplay video.
4. System Overview: Finite Automata Extraction
@@
In this section, we cover our proposed neurosymbolic approach to learn programs from gameplay video: Finite Automata Extraction (FAE). We give an overview of our pipeline in Figure 1. We begin by splitting the gameplay video data into RGB frames. We then train a self-supervised model based on the Marionette architecture (Smirnov et al., 2021) on these frames to extract a neural sprite dictionary, a bounded set of sprites. After training the model, we extract a learned symbolic representation of the frames. We then learn a program for each learned sprite to replicate their behaviour in the gameplay video. We learn each program by searching over Retro Coder, a novel domain-specific language.
4.1. Representation Extraction
@@
In this paper, we use entities to refer to unique 2D game objects. Each entity has a Sprite-Class, which dictates it’s appearence and behaviour. For example, Blinky, Pellets and Pac-Man would all be different Sprite-Classes. We require the set of all entities and the positional information of each entity in each frame for the later program learning step of FAE. We settled on a grid representation, where each grid is a matrix of symbolic values, where each value corresponds to a Sprite-Class. This enables us to handle collisions discretely. In order to obtain the grid representation, we chose to learn the sprite dictionary consisting of a maximum $l$ number of sprites using a variant of the Marionette architecture (Smirnov et al., 2021).
@@
We modified the original Marionette network to add an option to specify the background image, which is helpful for domains with a constant background. We changed the model to subtract the background image from both the input and output to ensure the model optimized only the reconstruction of the changing sprites. The original Marionette has multiple different loss components, including one for frame reconstruction. We added an additional multiplicative weight ($\gamma$) to this loss to further bias the model towards better reconstructions, which incentivizes learned sprite quality. We set $\gamma=10$ based on early experimentation.
@@
After training, for all the RGB frames $F_{1},...F_{n}$}, we extract their grid representations $G_{1},...G_{n}$}. Marionette employs a latent representation of frames in terms of grids of sprite embeddings. Thus we could extract these grids via passing the frames through the model and using the softmax weights from the model to determine the sprite present in each grid cell. This results in a $w\times h$ grid where each cell references a learned sprite in the sprite dictionary. As shown in Figure 2, the sprite dictionary consists of repeated sprites. For instance, index 0,1,2,7,8,13 and 14 all have sprites that resemble Pac-Man. To resolve this, we group similar looking sprites that imply similar behaviour together into same Sprite-Class. We then modify the $w\times h$ grid so that each cell in the grid refers to the Sprite-Class of the learned sprite.
Figure 2. Learned Sprite Dictionary
4.2. Retro Coder
@@
For this project we required a DSL that would support learning sprite behaviours. We thus defined our custom Domain Specific Language, Retro Coder. Retro Coder consists of condition functions (if) and action functions (then). Each if statement has one condition and one action. Since we designed FAE with 2D games in mind, we focused on common sprite behaviours and conditions for these behaviours. These consist of movement logic, entity transformation logic, and entity collision logic (Wardrip-Fruin, 2020). The condition functions are:
  • exists_in_map(sprite)exists\_in\_map(sprite) checks if an entity of sprite-class spritesprite exists in the frame
@@
  • $neighboring(sprite)$ checks if the entity of the program itself is next to an entity of sprite-class spritesprite,
@@
  • $neighbours(sprite1,sprite2)$ checks if an entity of sprite-class $sprite1$ is next to an entity of sprite-class $sprite2$. Since we are using a grid representation where one entity can exist at one location in a moment of time, we treat neighbouring sprites as collisions.
@@
  • follow_entity(sprite)follow\_entity(sprite) moves this entity toward the nearest entity of sprite-class spritesprite
  • follow_direction(direction)follow\_direction(direction) moves this entity in a direction (Up, down, left, right).
  • change_to_sprite(sprite)change\_to\_sprite(sprite) changes the entity’s sprite-class to spritesprite
  • follow_target_location(x,y)follow\_target\_location(x,y), which moves the entity towards to location $x,y$.
In the above we note that at times multiple entities of the same sprite-class. may appear on screen. For example, in Pac-man there are multiple pellets on screen at once.
4.3. Search Algorithm
@@
 for $batch\in(listofIndicesWhereSpriteExists,batchSize)$ do
@@
       $bestNeighbour\leftarrow program$;
@@
          if $D(batch,neighbour)<D(batch,bestNeighbour)$ then
@@
             $distanceDecreased\leftarrow True$;
@@
 $programs.\texttt{append}(program)$;
@@
Algorithm 1 Pseudocode for Search Algorithm
@@
This section covers the program synthesis step of FAE. The first step is that after we have our sprite-classes, we must categorize the sprite-classes into Learnable and Non-learnable sprites. This might be based on domain knowledge but can also be general at times. For example, the sprite-class that is controlled by the player would be Non-learnable as it does not have a ground truth program. We then learn a program for all of these hand-identified Learnable sprite-classes with our search algorithm, as seen in 1. We learn one program at a time, as this greatly reduces the search space. For a sprite-class $s$, we first scan $G_{1},...G_{n}$} to get a list of indices $I_{1},...I_{s}$} where $G_{I_{i}}$ contain $s$. We then go through the indices sequentially, for every $i$, compute the distance $D=\sum_{x=i}^{i+b}A.E(G_{I_{x+1}},G_{I_{x}})$ where $A.E$ is the Absolute Error and $b$ is the batch size. We set $b=3$ based on experimentation. We then search for neighbour programs to minimize $D$. The neighbour programs can either have an additional if-then statement, or remove an if-then statement from the current program. We consider all possible additions and all possible deletions in the $NeighbourPrograms()$ function.
5. Evaluation
Finite Automate Extraction allows us to process a single gameplay video of a 2D game and extract a learned set of sprites and a learned set of programs, one associated with each sprite. We evaluate our model from two perspectives: the quality of the learned programs, which assesses the backend of the model, and the quality of the world model as an approximation of a game, which assesses the front end. We test our model on two domains, Pac-Man and River Raid. Since this is an initial implementation, and thus primarily targeted at simple grid based 2D games, both of these meet the criteria while differing in terms of the types of games, with one focused on combat and the other on collecting.